# RefineStat: Efficient Exploration for Probabilistic Program Synthesis

### v1

@@ 3.2 Semantically-Constrained Probabilistic Program Generation
Figure 2: A constrained semantic decoding iteration in RefineStat
We formalize the generation of semantically valid probabilistic programs through iterative constrained sampling. Let $G=(\mathcal{N},\mathcal{T},\mathcal{P},S_{0})$ be a context-free grammar with nonterminal symbols $\mathcal{N}$, terminal symbols $\mathcal{T}$, production rules $\mathcal{P}$, and start symbol $S_{0}$. For a partial program $c\in L_{p}(G)$ with parse tree $\kappa$, we define validation functions that operate on parse tree nodes $n\in N(\kappa)$ within program context $\pi\in\Pi$:
$\Phi:N(\kappa)\times\Pi\mapsto\{0,1\}$
These functions are constructed as conjunctions of individual correctness checks:
@@
Here, the parse tree node $n$ is a program fragment which is semantically valid when all validation functions yield true.
Validity Predicates for Probabilistic Program Fragments. To ensure the correctness of synthesized probabilistic program fragments, we define three essential validation predicates. Each fragment $s\in S_{N}=\{\,s\in\mathcal{S}\mid s\text{ is derivable from nonterminal }N\in\mathcal{N}\,\}$ corresponds to a single statement in the program. Let $\mathcal{F}(s)$ denote the set of all probability distribution functions invoked within fragment $s$. The validation predicates are:
@@
    Parameter validity: $\phi_{3}(s,\Pi)=\prod_{f\in\mathcal{F}(s)}\mathbf{1}\{P(f)\subseteq P_{\mathrm{acc}}(f)\}$ confirms that operation parameters $P(f)$ adhere to the accepted specifications $P_{\mathrm{acc}}(f)$, essential for maintaining probabilistic semantics. i.e. we ensure that the provided parameters for any distribution are correct according to the specification of the distribution. For instance, as shown in the Figure 2, a use of parameter "std" was deemed invalid for the Normal distribution and it was resampled correctly as "sigma".
The complete validation predicate $\Phi(s,\Pi)=\phi_{1}(s,\Pi)\land\phi_{2}(s,\Pi)\land\phi_{3}(s,\Pi)$ ensures that generated fragments satisfy all requirements of the probabilistic programming language. Our generation approach leverages these semantics by maintaining a global symbol table $\Pi:\mathcal{A}\to\mathcal{M}$ mapping each alias $a\in\mathcal{A}$ to its module or namespace $m\in\mathcal{M}$.
Generation proceeds via local rejection sampling on $S_{N}$: we repeatedly sample $s\sim S_{N}$ until finding $s^{*}$ where $\Phi(s^{*},\Pi)=1$. This iterative process continues until a termination fragment is generated, ensuring every component in the final probabilistic program satisfies all semantic constraints. Note that our local rejection sampling is token-efficient and we backtrack and precisely resample the tokens that correspond to the violation of the constraints. The approach is particularly effective for languages like PyMC, where maintaining consistent probabilistic variable scopes and dependencies is critical. Combining syntactic constraints with semantic validation enables efficient exploration of the program space while ensuring the probabilistic soundness of generated models.
@@ 3.3 Program Validation and Guided Resampling
Algorithm 1 RefineStat Synthesis via $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$ with Constrained Decoding
@@
17:return $\arg\max_{M\in\mathcal{V}}\mathrm{elpd}(M)$
Building on the validation predicate $\Phi$, we now formalize our constrained generation and refinement methodology. We introduce the constrained-decoding operator $\mathcal{L}_{\mathrm{CD}}:\mathcal{C}\to\mathcal{B}$, which implements our validation-guided sampling. Here, a statement is an individual syntactic unit, and a code block $\mathcal{B}$ is a (possibly multi-statement) sequence of such statements. For any context $\mathcal{C}$, this operator returns a new code block $\mathcal{B}$ satisfying $\Phi(\mathcal{C}\|\mathcal{B})=1$, where $\|$ denotes sequential concatenation of blocks; concretely, for two blocks $A$ and $B$, $A\|B$ is the program text formed by appending all statements of $B$ immediately after those of $A$.
In particular, $\mathcal{D}\,\|\,\mathcal{P}\,\|\,\mathcal{L}$ denotes the full probabilistic program consisting first of the data block $\mathcal{D}$, then the prior block $\mathcal{P}$, and finally the likelihood block $\mathcal{L}$. During refinement we perform resampling via:
$\mathcal{L}\;\leftarrow\;L_{\textit{cd}}{}(\mathcal{D}\|\mathcal{P})\quad\text{(likelihood resampling)},$
$\mathcal{P}\;\leftarrow\;L_{\textit{cd}}{}(\mathcal{D})\quad\text{(prior resampling)},$
guaranteeing that each replaced block remains semantically valid.
Algorithm 1 synthesizes programs in two phases. First, it checks semantic correctness via $\Phi$, ensuring parseability, distribution validity, and parameter consistency. Second, it evaluates Bayesian diagnostics $d_{1},\dots,d_{L}$ on the full program $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$; at least $K$ thresholds $\{\tau_{j}\}$ must be met to accept a candidate. If diagnostics fail, we perform one of two resampling steps: (i) likelihood resampling replaces $\mathcal{L}$ under the data–prior context, addressing convergence or sampler-health issues; (ii) prior resampling replaces $\mathcal{P}$ under the data context, correcting prior-specification errors. We iterate until we collect $\beta$ valid programs or exhaust the rejection budget $R_{\max}$, and return the program maximizing $\widehat{\mathrm{elpd}}$.
@@ 4 Experimental Methodology
We use PyMC, a Python probabilistic programming library, to perform inference. We provide the same initial prompt while performing unconstrained generation and using RefineStat. We prompt the model by providing it with the dataset, the necessary library and the text query. The exact format of the prompt is provided in the Appendix C. To demonstrate RefineStat’s superiority, we allocate the Baseline a larger token budget than RefineStat, yet RefineStat still outperforms it despite using fewer tokens. Section 5.2 provides a detailed account of the experimental setup, including token allocation.
As defined in Definition 2 (Section 3.1), we assess model reliability using standard Bayesian diagnostics Vehtari et al. (2021, 2017); Gelman et al. (1995). We set the convergence threshold to $\alpha_{R}=1.05$ for split-$\widehat{R}$, allow $\mathrm{ESS}_{\mathrm{bulk}}\geq\beta_{\mathrm{bulk}}=400$, and adopt a relaxed cutoff $\beta_{\mathrm{tail}}=100$ for $\mathrm{ESS}_{\mathrm{tail}}$ to accommodate lower sampling efficiency in the tails. For leave-one-out validation, $\widehat{\mathrm{elpd}}(M)$ must be finite, with at least $1-\epsilon=0.8$ of data points having Pareto shape values below $L_{\mathrm{cd}}=0.7$. These thresholds are used consistently when computing the reliability score across all models. We use Standard unconstrained generation as our baseline. Further, based on preliminary experiments we have chosen $\beta$ to be 4, $\alpha$ to be 2, and $R_{\max}$ as 100 for all experimental purposes.
Datasets. We use five benchmark datasets from the Stan PosteriorDB Magnusson et al. (2024), mirroring their selection in prior research on automated statistical model generation Li et al. (2024):
@@
  • Dugongs Unit (b): This dataset provides measurements on the ages and lengths of 27 dugongs.
  • Surgical Unit (a): This dataset comprises records on the number of cardiac surgeries performed on infants, along with the associated failure rates.
@@
Experimental Setup. We run experiments on a 48-core Intel Xeon Silver 4214R CPU with 2 NVidia RTX A5000 GPUs. RefineStat is implemented using PyTorch Paszke et al. (2019), and Itergen library Ugare et al. (2024a) for refining the parser-guided LLM generation infrastructure. We run all experiments for 10 seeds to reduce result randomness, and use a temperature range of 0.2 to 0.4.
@@ 5 Experimental Results
In this section, we demonstrate how RefineStat enhances probabilistic programming model synthesis. We conduct a comprehensive evaluation on diverse datasets, comparing RefineStat against both an unconstrained baseline and leading language models (e.g., OpenAI o3) across a suite of diagnostic and performance metrics. Section 5.4 presents an ablation study that quantifies the contribution of each algorithmic component, metadata and contrasts RefineStat’s token consumption with that of the baseline model. Finally, we demonstrate that our method does not simply recall training data.
5.1 Improvement in Run Rate over Unconstrained and Syntax-driven Generation
To demonstrate the effectiveness of our method, we used identical prompts across our framework, the Standard baseline, Syncode Ugare et al. (2024c), and RefineStat. We compiled and ran each generated program to obtain the Run rates across different methods.
@@ 5.1 Improvement in Run Rate over Unconstrained and Syntax-driven Generation
Table 1 presents Run rates across temperature settings. Run rate is the fraction of programs that successfully produce the samples from the posterior distribution. The problems that do not run successfully include those with runtime errors such as (1) numerical errors, e.g., inf/nan, (2) sampling issues due to unlikely prior parameterization, (3) other sampling warnings, e.g., failed to initialize chains, (4) static compilation issues. RefineStat achieves success rates approximately 40 percentage points higher than the Standard baseline and 30 points higher than Syncode, demonstrating that our validation-guided approach substantially enhances code generation reliability by mitigating both syntactic and semantic error sources.These results show that RefineStat significantly enhances code generation reliability. Further, we categorized the failures by their root causes across different methods:
  • Syncode eliminated many basic syntactic mistakes but still suffered semantic errors, such as incorrect distribution parameter names, type mismatches, referring to deprecated API functions (e.g., calling pm.sample_prior from an earlier PyMC release), and inventing non-existent methods like pm.random_coefs.
  • RefineStat, in contrast, often produced models whose samplers failed to explore the correct posterior modes, leading to chains stuck in low-density regions or divergent transitions, failures stemming from the model definitions rather than our decoding framework; RefineStat reduced both syntactic and semantic errors and avoided sampler pathologies by enforcing grammar and parameter validity during decoding.
To evaluate our framework’s ability to produce higher quality probabilistic programs in terms of semantic correctness and diagnostic robustness, we ran two experimental conditions using the same prompt: the base language model alone, and RefineStat. Upon experimentation with RefineStat, we observed that on an average it consume almost twice the number of tokens consumed by the Baseline. Thus, to provide advantage to baseline models, we select the best program in terms of Bayesian Reliability Score, and $\widehat{\mathrm{elpd}}$ value among five different seeds for each dataset for comparison to a single run of RefineStat. We repeat this process five times for different seeds, to obtain the mean and standard deviation for all the metrics computed using these generated programs.
We then compare each generated program against the Bayesian Workflow Reliability score and the full suite of seven diagnostic criteria introduced earlier. To conserve space, we report five representative metrics drawn from these diagnostics. In our tables, the symbol ✗ indicates that no valid program was produced—i.e., the method failed to explore the search space sufficiently to yield a correct result.
DQ-7B failed on every dataset under the Standard Baseline, whereas it succeeded on all datasets when augmented with RefineStat. Except for one instance, RefineStat matches or exceeds the Standard Baseline in terms of the Bayesian Workflow Reliability score, and in some cases achieves up to twice the reliability. Moreover, RefineStat delivers substantially better performance on individual diagnostics, particularly divergences. For example, on the Surgical dataset with Meta-Llama, the Standard Baseline produces over 1000 divergences, while RefineStat produces none.
We observed that in cases where the Standard baseline attains a higher ELPD-LOO than RefineStat, closer inspection of the diagnostics exposes unreliable sampling. For instance, on the Meta-Llama GP task the Standard model achieves a superior ELPD score, but exhibits split-$\widehat{R}=4.13\gg 1$ and a low reliability score (3), indicating severe convergence issues. In contrast, RefineStat may report a marginally lower ELPD yet maintains $\widehat{R}\approx 1$ and a higher reliability score, reflecting trustworthy posterior estimates. Furthermore, RefineStat delivers markedly lower variability in key diagnostics such as $\widehat{R}$ and the number of divergent transitions demonstrating its consistency and robustness.
@@ 5.2 Comparison of Generated Program Quality to Unconstrained Baseline
Dataset Model Variant Reliab. Score $\uparrow$ $\widehat{R} \downarrow$ ESS Bulk $\uparrow$ Divergences $\downarrow$ Pareto $k$ $\downarrow$ ELPD LOO $\uparrow$
@@ 5.3 Comparison of Generated Program Quality to BoxLM
Table 3: Comparison with BoxLM Li et al. (2024), RefineStat using DQ-7B, and Expert values
Dataset Expert RefineStat BoxLM OpenAI-o3
w/ DQ-7B w/ GPT-4
Eight schools -30.70 -30.68 -30.42 -30.74
Dugongs 22.43 8.35 23.40 22.83
Peregrine -112.60 -114.29 -173.11 -133.29
Surgical -39.73 -46.51 -38.03 -38.73
GP -26.53 -23.39 -34.95
In this section, we compare the performance of our framework using the DQ-7B model (averaged over five runs) against three baselines: the Expert programs from PosteriorDB Magnusson et al. (2024), the BoxLM system introduced by Li et al. (2024), and programs generated by OpenAI o3. Since the code for BoxLM is not publicly available, we rely on the reported numbers from their paper for comparison. Additionally, because the specific dataset we use for the GP task is not included in BoxLM’s evaluation suite, we omit their result for that dataset in Table 3.
As shown in Table 3, our framework consistently outperforms both BoxLM and OpenAI o3 on the Peregrine dataset, and surpasses OpenAI o3 on the GP task, while matching the performance of expert-written programs in most cases. Across the remaining datasets, our approach performs comparably to other baselines, with the exception of Dugongs, where performance is slightly lower. These results demonstrate that RefineStat achieves performance comparable to, and in several cases better than, large language models like OpenAI o3 and multi-agent frameworks such as BoxLM.

### v2

@@ 3.1 Semantically-Constrained Probabilistic Program Generation
We formalize the generation of semantically valid probabilistic programs via iterative constrained sampling. Let $G=(\mathcal{N},\mathcal{T},\mathcal{P},S_{0})$ be a context-free grammar with nonterminal symbols $\mathcal{N}$, terminal symbols $\mathcal{T}$, production rules $\mathcal{P}$, and start symbol $S_{0}$.
For a partial program with parse tree $\kappa$, let $\mathcal{F}(\kappa)$ denote the set of program fragments, where each fragment $n\in\mathcal{F}(\kappa)$ is a rooted subtree of $\kappa$ corresponding to a single syntactic statement. Validation functions operate on a fragment $n$ within program context $\pi\in\Pi$: $\Phi:\mathcal{F}(\kappa)\times\Pi\to\{0,1\}.$ These functions are conjunctions of individual correctness checks:
@@
Each fragment thus corresponds to a single statement, possibly comprising multiple AST nodes.
Validity Predicates for Probabilistic Program Fragments. To ensure the correctness of synthesized probabilistic program fragments, we define three essential validation predicates. Let $\mathcal{F}(s)$ denote the set of all probability distribution functions invoked within fragment $s$. The validation predicates are:
@@
    Parameter validity: $\phi_{3}(s,\Pi)=\prod_{f\in\mathcal{F}(s)}\mathbf{1}\{P(f)\subseteq P_{\mathrm{acc}}(f)\}$ confirms that operation parameters $P(f)$ adhere to the accepted specifications $P_{\mathrm{acc}}(f)$, essential for maintaining probabilistic semantics. i.e. we ensure that the provided parameters for any distribution are correct according to the distribution’s specification. For instance, Figure 2 shows parameter "sd" was invalid and resampled correctly as "sigma".
  4. 4.
    Dependency validity: $\phi_{4}(s,\Pi)=\prod_{v\in\text{Vars}(s)}\mathbf{1}\{\text{all dependencies of }v\text{ are defined before use}\}$ ensures that random variables are declared and initialized before they are referenced.
  5. 5.
    Support validity: $\phi_{5}(s,\Pi)=\prod_{f\in\mathcal{F}(s)}\mathbf{1}\{P(f)\in\text{Supp}(f)\}$ confirms that parameter values fall within the distribution’s support (e.g., variance >0>0, probabilities in $[0,1]$).
  6. 6.
    Type validity: $\phi_{6}(s,\Pi)=\prod_{f\in\mathcal{F}(s)}\mathbf{1}\{\text{type}(P(f))\in T(f)\}$ checks that each parameter $P(f)$ has the expected type from the specification $T(f)$, e.g., ensuring numeric values for scale parameters, or integer values for counts.
Figure 2: Constrained decoding in RefineStat fixing a TypeError from using sd instead of sigma, as illustrated in Section 1.
The final predicate $\Phi(s,\Pi)=\bigwedge_{i=1}^{6}\phi_{i}(s,\Pi)$ ensures that generated fragments satisfy all requirements of the probabilistic programming language (1,4,6) and the Bayesian model (2,3,5). Our generation algorithm leverages these properties by maintaining a global symbol table $\Pi:\mathcal{A}\to\mathcal{M}$ mapping each alias $a\in\mathcal{A}$ to its module or namespace $m\in\mathcal{M}$.
Generation proceeds via local rejection sampling on $S_{N}$: we repeatedly sample $s\sim S_{N}$ until finding $s^{*}$ where $\Phi(s^{*},\Pi)=1$. This iterative process continues until a termination fragment is generated, ensuring every component in the final probabilistic program satisfies all semantic constraints. Our local rejection sampling is token-efficient and we backtrack and precisely resample the tokens that correspond to the violation of the constraints. The approach is particularly effective for languages like PyMC, where maintaining consistent probabilistic variable scopes and dependencies is critical. Combining syntactic constraints with semantic validation enables efficient exploration of the program space while ensuring the probabilistic soundness of generated models.
@@ 3.2 Program Validation and Guided Resampling
Algorithm 1 RefineStat Synthesis via $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$
@@
17:return $\arg\max\mathrm{elpd}(M)$ over all $M\in\mathcal{V}$}
Building on the validation predicate $\Phi$, we now formalize our constrained generation and refinement methodology. We introduce the constrained‐decoding operator $\mathcal{L}_{\mathrm{CD}}:\mathcal{C}\to\mathcal{B}$, which implements our validation‐guided sampling. Here, a statement is an individual syntactic unit, and a code block $\mathcal{B}$ is a (possibly multi‐statement) sequence of such statements. For any context $\mathcal{C}$, this operator returns a new code block $\mathcal{B}$ satisfying $\Phi(\mathcal{C}\|\mathcal{B})=1$, where $\|$ denotes sequential concatenation of blocks; concretely, for two blocks $A$ and $B$, $A\|B$ is the program text formed by appending all statements of $B$ immediately after those of $A$.
$\mathcal{D}\|\mathcal{P}\|\mathcal{L}$ denotes the full probabilistic program with data $\mathcal{D}$, prior $\mathcal{P}$, and likelihood $\mathcal{L}$. During refinement we perform resampling via two steps: $\mathcal{L}\leftarrow L_{\textit{cd}}(\mathcal{D}\|\mathcal{P})$ (likelihood resampling), $\mathcal{P}\leftarrow L_{\textit{cd}}(\mathcal{D})$ (prior resampling), guaranteeing replaced blocks remain semantically valid.
Before refinement begins, the data block $\mathcal{D}$ is taken directly from the user prompt, while the initial prior and likelihood blocks are generated via constrained decoding, i.e., $\mathcal{P}\leftarrow\mathcal{L}_{\mathrm{CD}}(\mathcal{D})$ and $\mathcal{L}\leftarrow\mathcal{L}_{\mathrm{CD}}(\mathcal{D}\mathbin{\|}\mathcal{P})$. Algorithm 1 synthesizes programs in two phases. First, it checks semantic correctness via $\Phi$, ensuring parseability, distribution validity, and parameter consistency. Second, it evaluates Bayesian diagnostics $d_{1},\dots,d_{L}$ on the full program $\mathcal{D}\|\mathcal{P}\|\mathcal{L}$; at least $K$ thresholds $\{\tau_{j}\}$ must be met to accept a candidate. As shown in Figure 1, if diagnostics fail, we perform one of two resampling steps to refine the program: (i) likelihood resampling replaces $\mathcal{L}$ under the data–prior context, addressing convergence or sampler‐health issues; (ii) prior resampling replaces $\mathcal{P}$ under the data context, correcting prior‐specification errors. We iterate until we collect $\beta$ valid programs or exhaust the budget $R_{\max}$, and return the program maximizing $\widehat{\mathrm{elpd}}$.
@@ 4 Experimental Methodology
We use PyMC, a Python probabilistic programming library, to perform inference. We provide the same initial prompt while performing unconstrained generation and using RefineStat. We prompt the model by providing it with the dataset, the necessary library and the text query. The exact format of the prompt is provided in the Appendix E. As stated in Definition 2, we assess model reliability using standard Bayesian diagnostics Vehtari et al. (2021; 2017); Gelman et al. (1995). Further details on hyperparameters and experimental setup are provided in Appendix F
Datasets. We use five benchmark datasets from Stan PosteriorDB Magnusson et al. (2024), mirroring the selection in prior research on automated statistical modeling  Li et al. (2024):
@@
  • Dugongs Unit : This dataset provides measurements on the ages and lengths of 27 dugongs.
  • Surgical Unit : This dataset comprises records on the number of cardiac surgeries performed on infants, along with the associated failure rates.
@@ 5 Experimental Results
5.1 Improved Run Rate over Unconstrained and Syntax-driven Generation
We conduct a comprehensive evaluation on diverse datasets, comparing RefineStat against both an unconstrained baseline and leading language models across a suite of diagnostic and performance metrics. We used identical prompts across our framework, the Standard baseline (unconstrained), syntactically-constrained tool Syncode Ugare et al. (2024c), and RefineStat.
Table 1 presents Run rates across different temperature settings. Run rate is the fraction of programs that successfully produce the samples from the posterior distribution. The problems that do not run successfully include those with runtime errors such as (1) numerical errors, e.g., inf/nan, (2) sampling issues due to unlikely prior parameterization, (3) other sampling warnings, e.g., failed to initialize chains, (4) static compilation issues.
@@ 5.1 Improved Run Rate over Unconstrained and Syntax-driven Generation
RefineStat achieves success rates approximately 40 percentage points higher than the Standard baseline and 30 points higher than Syncode, demonstrating that our validation-guided approach substantially enhances code generation reliability by mitigating both syntactic and
semantic error sources. These results show that RefineStat significantly enhances code generation reliability. The top root causes of failure are syntax errors, semantic errors, and sampler pathologies. We categorize these failures and provide a detailed discussion in Appendix G.1.
@@
[…]
@@ 5.2 Comparison of Generated Program Quality to Unconstrained Baseline
[…]
Except for one instance, RefineStat matches or exceeds the Standard Baseline in terms of the Bayesian Workflow Reliability score, and in some cases achieves up to twice the reliability. Moreover, RefineStat delivers substantially better performance on individual diagnostics, particularly divergences. Notably, DQ-7B, which failed on every dataset under the Standard Baseline, succeeded on all datasets when augmented with RefineStat. For example, on the Surgical dataset with Meta-Llama, the Standard Baseline produces over 1000 divergences, while RefineStat produces none.
We observed that in cases where the Standard baseline attains a higher ELPD‐LOO than RefineStat, closer inspection of the diagnostics exposes unreliable sampling. For instance, on the Meta‐Llama GP task the Standard model achieves a superior ELPD score, but exhibits split-$\widehat{R}=4.13\gg 1$ and a low reliability score (3), indicating severe convergence issues. In contrast, RefineStat may report a marginally lower ELPD yet maintains $\widehat{R}\approx 1$ and a higher reliability score, reflecting trustworthy posterior estimates. Furthermore, RefineStat delivers markedly lower variability in key diagnostics such as $\widehat{R}$ and the number of divergent transitions demonstrating its consistency and robustness. We further illustrate the types of structural changes introduced during refinement in Appendix D.2.
@@ 5.3 Comparison of Generated Program Quality to BoxLM
Table 3: Comparison of ELPD LOO scores with BoxLM Li et al. (2024), RefineStat using DQ-7B, and Expert values
Dataset Expert RefineStat w/ BoxLM OpenAI-o3
DQ-7B(mean $\pm$ std) w/ GPT-4 (mean $\pm$ std)
Eight schools -30.70 -30.68 $\pm$ 0.11 -30.42 -30.74 $\pm$ 0.07
Dugongs 22.43 8.35 $\pm$ 0.06 23.40 22.83 $\pm$ 8.12
Peregrine -112.60 -114.29 $\pm$ 2.76 -173.11 -133.29 $\pm$ 10.33
Surgical -39.73 -46.51 $\pm$ 3.49 -38.03 -38.73 $\pm$ 0.51
GP -26.53 -23.39 $\pm$ 1.14 -34.95 $\pm$ 13.28
We compare RefineStat’s performance using the DQ-7B model (averaged over five runs) against three baselines: the Expert stan programs from PosteriorDB Magnusson et al. (2024), the BoxLM system introduced by Li et al. (2024), and programs generated by OpenAI o3. Since the code for BoxLM is not publicly available, we rely on the reported numbers from their paper for comparison. Because the dataset we use for the GP task is not included in BoxLM’s evaluation suite, we omit their result for that dataset.
Table 3 presents the ELPD LOO scores, showing that our framework consistently outperforms both BoxLM and OpenAI o3 on the Peregrine dataset, and surpasses OpenAI o3 on the GP task, while matching the performance of expert-written programs in most cases. Across the remaining datasets, our approach performs comparably to other baselines, with the exception of Dugongs, where performance is slightly lower. These results demonstrate that RefineStat achieves performance comparable to, and in several cases better than, large language models like OpenAI o3 and multi-agent frameworks such as BoxLM.